# Revisiting the maximum mass limit of strange stars in quadratic curvature-matter coupled gravity

### v2

@@ I Introduction
Although there have been several attempts to model the interior of NS using different EoS, the proper explanation of physical properties of many stars (mass, radius, compactness etc.), estimated from recent observations, is not possible. The quark star hypothesis, where the compact star is hypothesised to be composed of Strange Quark Matter (SQM), has emerged as a prime candidate to describe the internal matter composition of such stars on the basis of observational results. Under the assumption of zero external pressure, Madsen [7] demonstrated that quark stars composed solely of $u$ and $d$ quarks are inherently unstable. However, the incorporation of $s$ quarks into the system significantly lowers the energy per baryon compared to the two-flavour configuration, thereby enhancing the stability. This necessitates the presence of $s$ quarks for a stable quark matter configuration. As a result, compact stars composed of $u$, $d$, and $s$ quarks are classified as Strange Stars (SS) or Strange Quark Stars (SQS) [8, 7]. Witten [9] further postulated that Strange Quark Matter (SQM), comprising approximately equal numbers of $u$, $d$, and $s$ quarks, may represent the true ground state of Quantum Chromodynamics (QCD), given that its energy per baryon is lower than that of the most stable atomic nucleus, 56Fe. This hypothesis lends strong theoretical support to the absolute stability of SQM. In the high-energy limit, QCD predicts asymptotic freedom, wherein nucleons dissociate into their fundamental constituents, forming a weakly interacting quark-gluon plasma (QGP) phase. Alford [10] proposed that the conditions in the dense and relatively cold cores of NS may be sufficient to trigger the formation of de-confined quark matter. Although the transition from hadronic matter to quark matter is not yet fully understood, various phenomenological approaches have been developed to model this process, viz., Nambu-Jona-Lasinio (NJL) model [11], Colour-Flavour-Locked (CFL) phase [12], Hyperonic EoS [13], kaon-meson condensate including hyperons [14], etc. Among these models, the MIT bag model EoS [15] has been one of the prime candidates to describe the interior of NS, through the thermodynamic behaviour of de-confined quark matter within a finite region, or ‘bag’. This bag represents the energy difference between the perturbative and non-perturbative QCD vacuum. Further, in MIT bag model, the quarks are considered to be degenerate Fermi gas, composed of $u$, $d$, $s$ quarks and electrons. Notably, several studies [16, 17, 18, 19] have been conducted to explain the properties of SS using the MIT bag model EoS.
@@
The correct extraction of the mass limit is only possible through observational evidence pertaining to the NS binary systems. Particularly, in the case of radio pulsars, the most accurate mass measurement provides a mass value of $1.35~\mathrm{M_{\odot}}$ [20]. In recent years, a significant volume of observational data from pulsars and gravitational wave (GW) events has been collected and analysed by various researchers. These analyses have led to precise estimations of several physical properties of compact astrophysical objects. Notably, using the Shapiro delay technique, the mass of the millisecond pulsar PSR J0740+6620 was measured to be $(2.14)^{+0.10}_{-0.09}~\mathrm{M{\odot}}$ [21], representing one of the most massive known NS. This measurement has intensified discussions regarding the upper mass limit of compact stars. Moreover, it has been suggested that certain compact objects in interacting binary systems could possess masses exceeding that of PSR J0740+6620 [21]. In particular, the gravitational wave event GW190814 revealed a secondary component with a mass in the range of $2.5$-$2.67~\mathrm{M{\odot}}$ at 90% confidence [22]. However, the exact nature of this object remains uncertain, it could either be an unusually massive NS, a SS or a low-mass black hole. If the former scenario holds true, it would necessitate theoretical advancements to extend the maximum mass to such high values. Developing such a framework is also crucial for probing the internal composition of matter at densities beyond nuclear saturation.
@@
However, achieving such high mass values and still maintaining the viability conditions in the context of General Relativity (GR) seems unlikely. Furthermore, the exact upper mass bound of NS/SS is still elusive, and day by day, the latest observational results show that we are in dire need of a new theory that can describe the recent astrophysical observations with greater accuracy. Therefore, we turn to the modified gravity approach. Modified gravity models, especially $f(R)$ [23, 24], $f(T)$ [25], $f(Q)$ [26], $f(G)$ [27], extended theories of gravity [28, 29, 27] etc. have demonstrated improved fits for known issues like cosmic acceleration [30], dark energy, and dark matter [31, 32, 33, 34, 35], and have been successfully applied to determining upper mass limits for compact stars [36, 37, 38, 39, 40, 41, 42, 43]. Among these, the $f(R)$ gravity model offers one of the most straightforward generalisations, in which the conventional Einstein-Hilbert action is substituted by a generic function of the Ricci scalar $R$ [44]. The $f(R)$ framework has provided significant insights into galactic kinematics and the accelerated expansion of the universe [45, 46]. Building on this, Harko et al. [47] extended the action further by incorporating the trace of the energy-momentum tensor $(T)$, leading to the $f(R,T)$ gravity theory. In their pioneering work, they interpreted the $T$ dependence as a potential signature of quantum effects or exotic imperfect fluids [48], and revealed distinctive features such as a non-vanishing covariant derivative of the energy-momentum tensor and an additional acceleration component arising from matter-curvature coupling.
@@
In this paper, we present a novel approach to redefine the maximum mass limit of SS, by considering the higher-order curvature corrections along with the gravity-matter coupling. This theory is effectively expressed as, $f(R,T)=R+\alpha R^{2}+2\beta T$ gravity, where $R$ and $T$ denote the Ricci scalar and trace of energy momentum tensor, respectively. Additionally, $\alpha$ and $\beta$ characterise the strength of higher-order curvature corrections and gravity-matter coupling. It must be noted that $f(R+\alpha R^{2})$ gravity is renowned for Starobinsky inflation [49], consistent with Planck 2018 data is cosmology [50]. This model is a more general framework of the modified gravity theories in comparison to the results obtained in Refs. [51, 52]. It is important to note that in the high-density, strong-gravity environment within NS/SS, higher-order curvature effects may arise naturally [53, 54, 55, 56]. Furthermore, to investigate the behaviour of matter under such extreme conditions and its influence on the EoS, and the resulting maximum mass limit, we have developed this new framework of an extended theory of gravity. Additionally, the new effective parameters arising from this theory can, potentially, remove some of the inconsistencies in the theoretical formalism of NS that are, in general, addressed through the use of exotic matter distribution. The main objective of this study is to incorporate quadratic curvature corrections and gravity-matter coupling into the TOV framework to derive the mass-radius relationship using the MIT bag model EoS. By systematically varying the parameters $\alpha$ and $\beta$, we aim to investigate the deviations from GR predictions, with a particular focus on identifying the emergence of strong gravity effects and their influence on the maximum mass limit of SS.
@@
In the present framework, we have coupled the matter sector to the higher-order curvature gravity. Now, a caveat is in order regarding the interpretation of the $2\beta T$ term. Although $T$ is a scalar built from matter, a term linear in $T$ in the gravitational action is not equivalent to the introduction of an independent propagating scalar field. The $R^{2}$ term produces a genuine scalar degree of freedom (the scalaron) after a Legendre/conformal transformation because it leads to a kinetic structure for that mode. In contrast, the $2\beta T$ term acts as a direct algebraic coupling between curvature and the matter trace and contains no kinetic term. Thus it modifies the geometry-matter coupling, and generally leads to $\nabla^{\mu}T_{\mu\nu}\neq 0$, but does not, by itself, introduce a new propagating scalar. If one desires a scalar-tensor representation where the matter trace is carried by an independent scalar field, that scalar must be introduced explicitly, for example, by promoting $T$ to a field $\chi$ with a kinetic term and potential. In this work we have maintained the phenomenological $2\beta T$ coupling and discussed its physical consequences for stellar structure. We have also examined limiting cases, and the dependence on the chosen matter Lagrangian. Previously, the studies of $f(R)$, or $f(R+\alpha R^{2})$ showed that in the high energy regime of compact stars, the inclusion of higher-order curvature provides a viable theoretical construct to study high mass compact stars. However, in presence of such high energy environment, the matter sector is bound to be influenced, and accounting for the matter contribution gives a complete foundation to study such extreme conditions of gravity. Keeping this in mind, we have constructed the present framework.
@@
Now, the main contents of the paper are organised as follows: Section II describes the mathematical formalism in the framework of $f(R+\alpha R^{2},T)$ gravity. Here, we develop the modified field equations for a static, spherically symmetric space-time and isotropic interior matter distribution. Section III deals with the formulation of TOV equations in this new context. The mass and pressure gradients are obtained from the modified field equations. The main results concerning the maximum mass and the associated radius are illustrated and tabulated in Section IV. In this section, we have obtained the particular ranges for the parameters that yield physically acceptable TOV solutions. Further, we have studied the stability of the model through the evaluation of adiabatic index in Section V. Finally, we summarise the key features, originating from this paper, in Section VI.
@@ II Formalism of quadratic curvature-matter coupled gravity
where we have considered the system of units, $G=1$ and $c=1$. In Eq. (1), the modified Ricci scalar is defined as, $\tilde{R}=R+\alpha R^{2}$. Consequently, $f(\tilde{R},T)$ is a function of Ricci scalar $R$, its quadratic correction $R^{2}$, governed by the coupling parameter $\alpha$, and the trace of energy momentum tensor $T$. Moreover, $\mathcal{L_{M}}$ is the Lagrangian associated with the matter sector. Now, varying Eq. (1) with respect to the fundamental metric tensor $g_{ij}$, we obtain:
@@
Here, $f_{R}=\frac{df(\tilde{R},T)}{dR}$, $f_{T}=\frac{df(\tilde{R},T)}{dT}$, $\Box\equiv~\frac{\partial_{i}(\sqrt{-g}g^{ij}\partial_{j})}{\sqrt{-g}}$ represents the d’Alembert operator, $\nabla_{i}$ denotes the covariant derivative associated with the Levi-Civita connection corresponding to the metric $g_{ij}$, and $\Theta_{ij}=g^{ij}\frac{\delta T_{ij}}{\delta g_{ij}}$ representing the variation of the energy-momentum tensor with respect to the metric. In the present formalism, considering the form, $f(\tilde{R},T)=R+\alpha R^{2}+2\beta T$ and refining Eq. (2), we obtain:
@@
where $\beta$ is the gravity-matter coupling parameter, $R_{ij}$ is the Ricci tensor and $G_{ij}$ represents the Einstein tensor. Notably, for $\alpha=0$ and $\beta=0$, we retain the field equations of GR. On the other hand, for $\alpha=0$ and $\beta\neq 0$, we obtain the $f(R,T)$ formalism presented by Harko et al. [47], whereas, for $\beta=0$ and $\alpha\neq 0$, we obtain the $f(R)$ formalism [57, 58].
@@ III Obtaining the Tolman-Oppenheimer-Volkoff equations
$=8\pi p+2p\beta-\beta\rho.$
Now, using the explicit result, $e^{-2\lambda}=1-\frac{2m(r)}{r}$ and the present form of $f(\tilde{R},T)$ model in Eqs. (7), (8) and (9) we obtain the modified TOV equation [2, 3] in the form:
@@ IV Maximum mass-radius using MIT Bag model EoS
In describing the interior of SS, MIT bag model [15] has been one of the most successful theories. Over time, several studies have been conducted that centred around the MIT bag model. Initially, the EoS contained a constant bag parameter $(B_{g})$ [60]. According to the predictions made by Madsen [7], the allowed range of the bag constant $B_{g}$ is constrained by the requirement of strange matter stability. The lower bound is set at $B_{g}^{1/4}=145~\mathrm{MeV}$, corresponding to $B_{g}=57.55~\mathrm{MeV/fm^{3}}$, below which strange quark matter becomes unstable. In scenarios involving only $u$ and $d$ quarks, the system may not be stable energetically, as also indicated in Ref. [7]. The inclusion of the $s$-quark lowers the energy per baryon, thereby enhancing the stability of the system. For two-flavour (non-strange) quark matter to be stable relative to nuclear matter, its energy per baryon must remain below the neutron mass, $939.6~\mathrm{MeV}$ [7] at zero external pressure. This condition sets the minimum value of $B_{g}$ necessary to prevent atomic nuclei from decaying into non-strange quark matter. On the other hand, the upper bound on $B_{g}$ arises from the requirement that strange quark matter be stable compared to iron nuclei. This bound corresponds to $(B_{g})^{1/4}_{\mathrm{{max}}}=162.8~\mathrm{MeV}$ or equivalently $(B_{g})_{\mathrm{{max}}}=91.54~\mathrm{MeV/fm^{3}}$ [7]. When considering stability relative to a neutron under zero external pressure, the upper limit shifts slightly to $(B_{g})^{1/4}_{\mathrm{{max}}}=164.4~\mathrm{MeV}$, yielding $(B_{g})_{\mathrm{{max}}}=95.11~\mathrm{MeV/fm^{3}}$ [61, 7].
In this section, we have solved the modified TOV equations using the MIT EoS for the upper and lower bounds of the bag parameter, $B_{g}$. Notably, the choice of $\alpha$ and $\beta$ is crucial here. In higher-order curvature-matter coupled gravity, the appearance of ghost field is a problematic scenario. To address this situation, we have ensured a positive kinetic term by imposing the condition $f_{RR}\geq 0$. In the present model, $f_{RR}\geq 0$ leads to $\alpha>0$. Further, following the previous studies [62, 63], we have noted that $\beta$ can take values across both positive and negative ranges. Hence, we have considered systematic variations in $\alpha$ and $\beta$ to obtain the maximum mass-radius relation. Notably, the ranges of $\beta$ and $\alpha$ are chosen carefully so that the solution of the TOV equation remains well-behaved. Based on this setup, the resulting findings are presented and discussed below:
@@
    • From Fig. 1, we note that for a negative choice of $\beta$, the maximum mass-radius increases with increasing $\alpha$. This may be related to the fact that negative gravity-matter coupling indicates a weakly coupled system that allows for a larger configuration. Further, with increasing curvature corrections, the stellar structure may counter gravity more strongly and withstand the gravitational collapse. Consequently, the mass and the associated radius increase. However, from Table 1(a), we have noted that for $\beta=-0.5$, we can correctly extract the maximum mass and radius up to $\alpha=4.7$, beyond which TOV equations cannot be solved properly.
@@
V Stability Analysis through radial variation of adiabatic index

### v3

# Maximum mass limit of strange stars in quadratic curvature-matter coupled gravity

@@ I Introduction
Although there have been several attempts to model the interior of NS using different EoS, the proper explanation of physical properties of many stars (mass, radius, compactness etc.), estimated from recent observations, is not possible. The quark star hypothesis, where the compact star is hypothesised to be composed of Strange Quark Matter (SQM), has emerged as a prime candidate to describe the internal matter composition of such stars on the basis of observational results. Under the assumption of zero external pressure, Madsen [7] demonstrated that quark stars composed solely of $u$ and $d$ quarks are inherently unstable. However, the incorporation of $s$ quarks into the system significantly lowers the energy per baryon compared to the two-flavour configuration, thereby enhancing the stability. This necessitates the presence of $s$ quarks for a stable quark matter configuration. As a result, compact stars composed of $u$, $d$, and $s$ quarks are classified as Strange Stars (SS) or Strange Quark Stars (SQS) [8, 7]. Witten [9] further postulated that Strange Quark Matter (SQM), comprising approximately equal numbers of $u$, $d$, and $s$ quarks, may represent the true ground state of Quantum Chromodynamics (QCD), given that its energy per baryon is lower than that of the most stable atomic nucleus, 56Fe. This hypothesis lends strong theoretical support to the absolute stability of SQM. In the high-energy limit, QCD predicts asymptotic freedom, wherein nucleons dissociate into their fundamental constituents, forming a weakly interacting quark-gluon plasma (QGP) phase. Alford [10] proposed that the conditions in the dense and relatively cold cores of NS may be sufficient to trigger the formation of de-confined quark matter. Although the transition from hadronic matter to quark matter is not yet fully understood, various phenomenological approaches have been developed to model this process, viz., MIT bag model EoS [11], Nambu-Jona-Lasinio (NJL) model [12], Colour-Flavour-Locked (CFL) phase [13], Hyperonic EoS [14], kaon-meson condensate including hyperons [15], etc. Among these models, the MIT bag model EoS [11] has been one of the prime candidates to describe the interior of SQS, through the thermodynamic behaviour of de-confined quark matter within a finite region, or ‘bag’. This bag represents the energy difference between the perturbative and non-perturbative QCD vacuum. Further, in MIT bag model, the quarks are considered to be degenerate Fermi gas, composed of $u$, $d$, $s$ quarks and electrons. Notably, several studies [16, 17, 18, 19] have been conducted to explain the properties of SS using the MIT bag model EoS.
@@
The correct extraction of the mass limit is only possible through observational evidence pertaining to the NS binary systems. Particularly, in the case of radio pulsars, the most accurate mass measurement provides a mass value of $1.35~\mathrm{M_{\odot}}$ [20]. In recent years, a significant volume of observational data from pulsars and gravitational wave (GW) events has been collected and analysed by various researchers. These analyses have led to precise estimations of several physical properties of compact astrophysical objects. Notably, using the Shapiro delay technique, the mass of the millisecond pulsar PSR J0740+6620 was measured to be $(2.14)^{+0.10}_{-0.09}~\mathrm{M{\odot}}$ [21], representing one of the most massive known pulsar. This measurement has intensified discussions regarding the upper mass limit of compact stars. Moreover, it has been suggested that certain compact objects in interacting binary systems could possess masses exceeding that of PSR J0740+6620 [21]. In particular, the gravitational wave event GW190814 revealed a secondary component with a mass in the range of $2.5$-$2.67~\mathrm{M{\odot}}$ at 90% confidence [22]. However, the exact nature of this object remains uncertain, it could either be an unusually massive NS, a SS or a low-mass black hole. If the former scenario holds true, it would necessitate theoretical advancements to extend the maximum mass to such high values. Developing such a framework is also crucial for probing the internal composition of matter at densities beyond nuclear saturation.
@@
However, achieving such high mass values and still maintaining the viability conditions in the context of General Relativity (GR) seems unlikely. Furthermore, the exact upper mass bound of NS/SS is still elusive, and day by day, the latest observational results show that we are in dire need of new theory that can describe the recent astrophysical observations with greater accuracy. Therefore, we turn to the modified gravity approach. Modified gravity models, especially $f(R)$ [23, 24], $f(T)$ [25], $f(Q)$ [26], $f(G)$ [27], extended theories of gravity [28, 29, 27] etc. have demonstrated improved fits for known issues like cosmic acceleration [30], dark energy, and dark matter [31, 32, 33, 34, 35], and have been successfully applied to determining upper mass limits for compact stars [36, 37, 38, 39, 40, 41, 42, 43]. Among these, the $f(R)$ gravity model offers one of the most straightforward generalisations, in which the conventional Einstein-Hilbert action is substituted by a generic function of the Ricci scalar $R$ [44]. The $f(R)$ framework has provided significant insights into galactic kinematics and the accelerated expansion of the universe [45, 46]. Building on this, Harko et al. [47] extended the action further by incorporating the trace of the energy-momentum tensor $(T)$, leading to the $f(R,T)$ gravity theory. In their pioneering work, they interpreted the $T$ dependence as a potential signature of quantum effects or exotic imperfect fluids [48], and revealed distinctive features such as a non-vanishing covariant derivative of the energy-momentum tensor and an additional acceleration component arising from matter-curvature coupling.
@@
In this paper, we present a novel approach to redefine the maximum mass limit of SS, by considering the higher-order curvature corrections along with the gravity-matter coupling. This theory is effectively expressed as, $f(R,T)=R+\alpha R^{2}+2\beta T$ gravity, where $R$ and $T$ denote the Ricci scalar and trace of energy momentum tensor, respectively. Additionally, $\alpha$ and $\beta$ characterise the strength of higher-order curvature corrections and gravity-matter coupling. It must be noted that $f(R+\alpha R^{2})$ gravity is renowned for Starobinsky inflation [49], consistent with Planck 2018 data in cosmology [50]. This model is a more general framework of the modified gravity theories in comparison to the results obtained in Refs. [51, 52]. It is important to note that in the high-density, strong-gravity environment within NS/SS, higher-order curvature effects may arise naturally [53, 54, 55, 56]. Furthermore, to investigate the behaviour of matter under such extreme conditions and its influence on the EoS, and the resulting maximum mass limit, we have developed this new framework of an extended theory of gravity. Additionally, the new effective parameters arising from this theory can, potentially, remove some of the inconsistencies in the theoretical formalism of compact stars that are, in general, addressed through the use of exotic matter distribution. The main objective of this study is to incorporate quadratic curvature corrections and gravity-matter coupling into the TOV framework to derive the mass-radius relationship using the MIT bag model EoS. By systematically varying the parameters $\alpha$ and $\beta$, we aim to investigate the deviations from GR predictions, with a particular focus on identifying the emergence of strong gravity effects and their influence on the maximum mass limit of SS.
@@
In the present framework, we couple the matter sector to the higher-order curvature gravity. A caveat is in order regarding the interpretation of the $2\beta T$ term. Although $T$ is a scalar built from matter, a term linear in $T$ in the gravitational action is not equivalent to the introduction of an independent propagating scalar field. The $R^{2}$ term produces a genuine scalar degree of freedom (the scalaron) after a Legendre/conformal transformation because it leads to a kinetic structure for that mode. In contrast, the $2\beta T$ term acts as a direct algebraic coupling between curvature and the matter trace and contains no kinetic term. Thus it modifies the geometry-matter coupling, and generally leads to $\nabla^{\mu}T_{\mu\nu}\neq 0$, but does not, by itself, introduce a new propagating scalar. If one desires a scalar-tensor representation where the matter trace is carried by an independent scalar field, that scalar must be introduced explicitly, for example, by promoting $T$ to a field $\chi$ with a kinetic term and potential. In this work we maintain the phenomenological $2\beta T$ coupling and discussed its physical consequences for stellar structure. We also examine limiting cases, and the dependence on the chosen matter Lagrangian. Previously, the studies of $f(R)$, or $f(R+\alpha R^{2})$ showed that in the high energy regime of compact stars, the inclusion of higher-order curvature provides a viable theoretical construct to study high mass compact stars. However, in presence of such high energy environment, the matter sector is bound to be influenced, and accounting for the matter contribution gives a complete foundation to study such extreme conditions of gravity. Keeping this in mind, we construct the present framework.
@@
The main contents of the paper are organised as follows: Section II describes the mathematical formalism in the framework of $f(R+\alpha R^{2},T)$ gravity. Here, we develop the modified field equations for a static, spherically symmetric space-time and isotropic interior matter distribution. Section III deals with the formulation of TOV equations in this new context. The mass and pressure gradients are obtained from the modified field equations. The main results concerning the maximum mass and the associated radius are illustrated and tabulated in Section IV. In this section, we obtain the particular ranges for the parameters that yield physically acceptable TOV solutions. Further, we study the stability of the model through the evaluation of adiabatic index and Harrison-Zel’dovich-Novikov condition in Section V. The possible tests that may lead to strange star detection, thereby validating the present framework, are addressed in Section VI. Finally, we summarise the key features, originating from this paper, in Section VI.
@@ II Formalism of quadratic curvature-matter coupled gravity
where we have considered the system of units, $G=1$ and $c=1$. In Eq. (1), the modified Ricci scalar is defined as, $\tilde{R}=R+\alpha R^{2}$, where $\alpha$ is the curvature correction parameter. In the present analysis, we consider the following form: $f(\tilde{R},T)=R+\alpha R^{2}+2\beta T$, which explicitly includes the quadratic curvature modifications along with the gravity-matter coupling, through the trace of energy-momentum tensor $T$, governed by the coupling parameter $\beta$. Moreover, $\mathcal{L_{M}}$ is the Lagrangian associated with the matter sector. Notably, this form of $f(\tilde{R},T)$ provides the following advantages and motivations:
  • The $R^{2}$ correction arises naturally in semi-classical gravity as a one-loop quantum effect, and it supports a successful high-curvature inflationary phase (Starobinsky inflation).
  • In the cosmological aspect, these type of models may dynamically support self-consistent de Sitter-like solutions during large $R$.
  • In such matter-gravity coupled theories, the viability conditions $f_{R}>0$ and $f_{RR}>0$ demand $\alpha>0$, which will avoid ghost modes and the Dolgov-Kawasaki instability [57].
  • In such extended $f(R,T)$ model, a small $\alpha$ yields a sufficiently massive scalaron, recovering GR behaviour in the low-curvature regime and satisfying Solar-System bounds.
  • The linear $T$-dependence enables matter-geometry coupling governed by $\beta$, offering phenomenological flexibility in strong-gravity environments such as compact stars.
  • This model may predict stable, causal interior solutions with realistic EoS when $\alpha>0$ and $|\beta|$ is small.
  • The theory smoothly recovers GR in the limit $\alpha\rightarrow 0$ and $\beta\rightarrow 0$, indicating that the present framework is a minimal and well-motivated modification of Einstein gravity.
Now, varying Eq. (1) with respect to the fundamental metric tensor $g_{ij}$, we obtain:
@@
Here, $f_{R}=\frac{df(\tilde{R},T)}{dR}$, $f_{T}=\frac{df(\tilde{R},T)}{dT}$, $\Box\equiv~\frac{\partial_{i}(\sqrt{-g}g^{ij}\partial_{j})}{\sqrt{-g}}$ represents the d’Alembert operator, $\nabla_{i}$ denotes the covariant derivative associated with the Levi-Civita connection corresponding to the metric $g_{ij}$, and $\Theta_{ij}=g^{ij}\Big(\frac{\delta T_{ij}}{\delta g_{ij}}\Big)$ representing the variation of the energy-momentum tensor with respect to the metric. Now considering the form, $f(\tilde{R},T)=R+\alpha R^{2}+2\beta T$ and refining Eq. (2), we obtain:
@@
where $R_{ij}$ is the Ricci tensor and $G_{ij}$ represents the Einstein tensor. Notably, for $\alpha=0$ and $\beta\neq 0$, we obtain the $f(R,T)$ formalism presented by Harko et al. [47], whereas, for $\beta=0$ and $\alpha\neq 0$, we obtain the $f(R)$ formalism [58, 59].
@@ III Obtaining the Tolman-Oppenheimer-Volkoff equations
$=8\pi p+3p\beta-\beta\rho.$
Now, using the explicit result, $e^{-2\lambda}=1-\frac{2m(r)}{r}$ and the present form of $f(\tilde{R},T)$ model in Eqs. (7), (8) and (9), we obtain the modified TOV equation [2, 3] in the form:
@@ IV Maximum mass-radius using MIT Bag model EoS
In describing the interior of SS, MIT bag model [11] has been one of the most successful theories. Over time, several studies have been conducted that centred around the MIT bag model. Initially, the EoS contained a constant bag parameter $(B_{g})$ [61]. According to the predictions made by Madsen [7], the allowed range of the bag constant $B_{g}$ is constrained by the requirement of strange matter stability. The lower bound is set at $B_{g}^{1/4}=145~\mathrm{MeV}$, corresponding to $B_{g}=57.55~\mathrm{MeV/fm^{3}}$, below which non-interacting strange quark matter becomes unstable. In scenarios involving only $u$ and $d$ quarks, the system may not be stable energetically, as also indicated in Ref. [7]. The inclusion of the $s$-quark lowers the energy per baryon, thereby enhancing the stability of the system. For two-flavour (non-strange) quark matter to be stable relative to nuclear matter, its energy per baryon must remain below the neutron mass, $939.6~\mathrm{MeV}$ [7] at zero external pressure. This condition sets the minimum value of $B_{g}$ necessary to prevent atomic nuclei from decaying into non-strange quark matter. On the other hand, the upper bound on $B_{g}$ arises from the requirement that strange quark matter be stable compared to iron nuclei. This bound corresponds to $(B_{g})^{1/4}_{\mathrm{{max}}}=162.8~\mathrm{MeV}$ or equivalently $(B_{g})_{\mathrm{{max}}}=91.54~\mathrm{MeV/fm^{3}}$ [7]. When considering stability relative to a neutron under zero external pressure, the upper limit shifts slightly to $(B_{g})^{1/4}_{\mathrm{{max}}}=164.4~\mathrm{MeV}$, yielding $(B_{g})_{\mathrm{{max}}}=95.11~\mathrm{MeV/fm^{3}}$ [62, 7].
In this section, we have solved the modified TOV equations using the MIT EoS for the upper and lower bounds of the bag parameter, $B_{g}$. Notably, the choice of $\alpha$ and $\beta$ is crucial here. In higher-order curvature-matter coupled gravity, the appearance of ghost field is a problematic scenario. To address this situation, we have ensured a positive kinetic term by imposing the condition $f_{RR}\geq 0$. In the present model, $f_{RR}\geq 0$ leads to $\alpha>0$. Further, following the previous studies [63, 64], we have noted that $\beta$ can take values across both positive and negative ranges. Hence, we have considered systematic variations in $\alpha$ and $\beta$ to obtain the maximum mass-radius relation. Notably, the ranges of $\alpha$ and $\beta$ are chosen carefully so that the solution of the TOV equation remains well-behaved. Based on this setup, the resulting findings are presented and discussed below:
@@
    • From Fig. 1, we note that for a negative choice of $\beta$, the maximum mass-radius increases with increasing $\alpha$. This may be related to the fact that negative gravity-matter coupling indicates a weakly coupled system that allows for a larger configuration. Further, with increasing curvature corrections, the stellar structure may counter gravity more strongly and withstand the gravitational collapse. Consequently, the mass and the associated radius increase. However, from Table 1(a), we have noted that for $\beta=-0.5$, we can extract the maximum mass and radius up to $\alpha=4.7$, beyond which TOV equations cannot be solved properly.
@@
V Stability Analysis
The validity of any newly proposed framework relies on the stability analysis. In the present context, since we have used a particular barotropic EoS, we have studied the variation of adiabatic index to substantiate this aspect. Further, the proposed analysis revolves around the determination of maximum mass-radius in this novel framework, hence we have studied the Harrison-Zel’dovich-Novikov criterion [70] to assess the stability of the model. Notably, we have selected the following range of parameters in for this analysis: $B_{g}=57.55$ and $95.11~\mathrm{MeV/fm^{3}}$, $\beta=0.5$, and $\alpha=0,~2.5$ and $5$.
V.1 Radial variation of adiabatic index